已知事件 AA 和 BB 满足 P(B)=13\mathrm{P}(B) = \frac{1}{3},P(A∣B)=25\mathrm{P}(A|B) = \frac{2}{5} 且 P(A∪B)=35\mathrm{P}(A \cup B) = \frac{3}{5}。

  1. (i)

    求 P(A)\mathrm{P}(A),并由此判断事件 AA 和 BB 是否独立。

  2. (ii)

    求 P(A′∩B)\mathrm{P}(A' \cap B) 的精确值。

  3. (iii)

    对于第三个事件 CC,已知 P(C)=12\mathrm{P}(C) = \frac{1}{2} 且 AA 和 CC 互斥。

    准确求出 P(B′∩C)\mathrm{P}(B' \cap C) 的最大和最小可能值。

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

9(i)

1. Calculate the probability of the intersection of A and B using the conditional probability formula.

P(A∩B)=P(A∣B)×P(B)=25×13=215\mathrm{P}(A \cap B) = \mathrm{P}(A|B) \times \mathrm{P}(B) = \frac{2}{5} \times \frac{1}{3} = \frac{2}{15}

2. Find $\mathrm{P}(A)$ using the addition rule of probability.

P(A∪B)=P(A)+P(B)−P(A∩B)\mathrm{P}(A \cup B) = \mathrm{P}(A) + \mathrm{P}(B) - \mathrm{P}(A \cap B)
35=P(A)+13−215\frac{3}{5} = \mathrm{P}(A) + \frac{1}{3} - \frac{2}{15}
P(A)=35−13+215=9−5+215=615=25\mathrm{P}(A) = \frac{3}{5} - \frac{1}{3} + \frac{2}{15} = \frac{9 - 5 + 2}{15} = \frac{6}{15} = \frac{2}{5}

3. Determine if events A and B are independent by checking the mathematical conditions for independence.

Since P(A∣B)=25\mathrm{P}(A|B) = \frac{2}{5} and P(A)=25\mathrm{P}(A) = \frac{2}{5}, we observe that P(A∣B)=P(A)\mathrm{P}(A|B) = \mathrm{P}(A).

Alternatively, checking the product rule:
P(A)×P(B)=25×13=215\mathrm{P}(A) \times \mathrm{P}(B) = \frac{2}{5} \times \frac{1}{3} = \frac{2}{15}
Since P(A)×P(B)=P(A∩B)\mathrm{P}(A) \times \mathrm{P}(B) = \mathrm{P}(A \cap B), events AA and BB are independent.

答案

P(A)=25\mathrm{P}(A) = \frac{2}{5}, events AA and BB are independent

9(ii)

1. Calculate the probability of the intersection of $A$ and $B$.

P(A∩B)=P(A∣B)P(B)=25×13=215\mathrm{P}(A \cap B) = \mathrm{P}(A|B)\mathrm{P}(B) = \frac{2}{5} \times \frac{1}{3} = \frac{2}{15}

2. Calculate $\mathrm{P}(A' \cap B)$ using the rule for the difference of events.

P(A′∩B)=P(B)−P(A∩B)=13−215=515−215=315=15\mathrm{P}(A' \cap B) = \mathrm{P}(B) - \mathrm{P}(A \cap B) = \frac{1}{3} - \frac{2}{15} = \frac{5}{15} - \frac{2}{15} = \frac{3}{15} = \frac{1}{5}

答案

15\frac{1}{5}

9(iii)

1. Define variables for the probabilities of the relevant regions within $A'$ to formulate equations.

Since AA and CC are mutually exclusive, P(A∩C)=0\mathrm{P}(A \cap C) = 0. This means event CC occurs entirely within A′A'.
Let y=P(A′∩B′∩C)y = \mathrm{P}(A' \cap B' \cap C). Since P(A∩B′∩C)=0\mathrm{P}(A \cap B' \cap C) = 0, y=P(B′∩C)y = \mathrm{P}(B' \cap C).
Let x=P(A′∩B∩C)x = \mathrm{P}(A' \cap B \cap C).
Event CC is partitioned into its intersections with BB and B′B', so:
x+y=P(C)=12x + y = \mathrm{P}(C) = \frac{1}{2}

2. Use the probability of $A' \cap B$ and the non-negativity of probabilities to bound $x$ and find the least value of $y$.

From part (ii), we have P(A′∩B)=15\mathrm{P}(A' \cap B) = \frac{1}{5}.
The region A′∩BA' \cap B is partitioned by CC and C′C'. Thus, the probability of the part outside CC is:
P(A′∩B∩C′)=15−x\mathrm{P}(A' \cap B \cap C') = \frac{1}{5} - x
Since probabilities cannot be negative, 15−x≥0\frac{1}{5} - x \ge 0, which gives x≤15x \le \frac{1}{5}.
Substitute x=12−yx = \frac{1}{2} - y:
12−y≤15  ⟹  y≥12−15=310\frac{1}{2} - y \le \frac{1}{5} \implies y \ge \frac{1}{2} - \frac{1}{5} = \frac{3}{10}
The least possible value of P(B′∩C)\mathrm{P}(B' \cap C) is 310\frac{3}{10}.

3. Use the total probability of $A'$ to bound $y$ and find its greatest value.

From part (i), we established P(A)=25\mathrm{P}(A) = \frac{2}{5}, so P(A′)=1−25=35\mathrm{P}(A') = 1 - \frac{2}{5} = \frac{3}{5}.
Let z=P(A′∩B′∩C′)z = \mathrm{P}(A' \cap B' \cap C').
The event A′A' comprises four mutually exclusive regions, so their probabilities must sum to P(A′)\mathrm{P}(A'):
P(A′∩B∩C)+P(A′∩B∩C′)+P(A′∩B′∩C)+P(A′∩B′∩C′)=P(A′)\mathrm{P}(A' \cap B \cap C) + \mathrm{P}(A' \cap B \cap C') + \mathrm{P}(A' \cap B' \cap C) + \mathrm{P}(A' \cap B' \cap C') = \mathrm{P}(A')
x+(15−x)+y+z=35x + \left(\frac{1}{5} - x\right) + y + z = \frac{3}{5}
15+y+z=35  ⟹  y+z=25\frac{1}{5} + y + z = \frac{3}{5} \implies y + z = \frac{2}{5}
Since z≥0z \ge 0, we must have y≤25y \le \frac{2}{5}.
The greatest possible value of P(B′∩C)\mathrm{P}(B' \cap C) is 25\frac{2}{5}.

答案

Greatest value =25= \frac{2}{5}, Least value =310= \frac{3}{10}

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1凯(Kai)和里奥(Leo)两兄弟是定期训练的双项全能(跑步和骑行)运动员。在每一天中,凯骑行的概率为 34\frac{3}{4},跑步的概率为 35\frac{3}{5},两者都进行的概率为 pp。 平均而言,里奥在一周的 7 天中有…
  2. 2一对同胞 Abbey 和 Betty 还有其他六个朋友一起去参加书展。他们八个人随机排成一队等候一位著名作家的签名。求满足以下条件的概率: Abbey 排在队列第一位且 Betty 排在队列第六位, Abbey 排在队列第一位或 Betty…
  3. 3已知某种特定疾病的感染率约为万分之一(每 10 000 人中约有 1 人感染)。针对该疾病的一项检测准确率为 99.9%,即检测结果与真实情况相符的概率为 99.9%。求在已知某人检测结果呈阳性的条件下,此人确实患有该疾病的概率。 研究人员…
  4. 4对于事件 AA 和 BB,已知 P(A)=0.55\mathrm{P}(A) = 0.55,P(B)=0.65\mathrm{P}(B) = 0.65 且 P(A∪B)=0.93\mathrm{P}(A \cup B) = 0.93。 求 P(A∩B)\mathrm{P}(A \cap B)…
  5. 5对于事件 AA 和 BB,已知 P(A′)=0.3\mathrm{P}(A') = 0.3 且 P(B)=0.55\mathrm{P}(B) = 0.55。 已知事件 AA 和 BB 相互独立,求 P(A∩B)\mathrm{P}(A \cap B)。 已知事件 AA…